Consider an intrinsic semiconductor crystal at room temperature, where kBT is 0.025 eV. The probability of a state close to the valence-band edge being occupied by a hole is 1.0 x 10-5. Calculate the band gap.


1
Expert's answer
2021-05-23T16:36:12-0400

"k_BT=0.025 \space eV"

"f(E_c)=1\times10^{-5}"

"f(E_c)=e^{-E_g\/2k_BT}"

Taking natural log both sides

"\ln(f(E_c))=\dfrac{-E_g}{2k_BT}"

"E_g=-2k_BT\ln(f(E_c))"

"E_g=0.575\space eV"

Bandgap = 0.575 eV